### v1

# Cross-domain Multi-modal Few-shot Object Detection via Rich Text

Abstract
Cross-modal feature extraction and integration have led to steady performance improvements in few-shot learning tasks due to generating richer features. However, existing multi-modal object detection (MM-OD) methods degrade when facing significant domain-shift and are sample insufficient. We hypothesize that rich text information could more effectively help the model to build a knowledge relationship between the vision instance and its language description and can help mitigate domain shift. Specifically, we study the Cross-Domain few-shot generalization of MM-OD (CDMM-FSOD) and propose a meta-learning based multi-modal few-shot object detection method that utilizes rich text semantic information as an auxiliary modality to achieve domain adaptation in the context of FSOD. Our proposed network contains (i) a multi-modal feature aggregation module that aligns the vision and language support feature embeddings and (ii) a rich text semantic rectify module that utilizes bidirectional text feature generation to reinforce multi-modal feature alignment and thus to enhance the model’s language understanding capability. We evaluate our model on common standard cross-domain object detection datasets and demonstrate that our approach considerably outperforms existing FSOD methods. Our implementation is publicly available: https://github.com/zshanggu/CDMM
Keywords:
Few-shot Object Detection Multi-modal Cross-domain
1 Introduction
The goal in few-shot object detection (FSOD) is to train a model that can generalize to detecting objects for which we only have a few labeled examples [56, 46]. This task has applications in domains where encountering rare or less common objects is frequent, such as autonomous driving. A typical FSOD training method follows a two-fold training procedure. First, a model is pre-trained on a number of base classes for which we have abundant data and then the model is fine-tuned on a set of novel classes with few annotated samples. In the classic setting, the base and the novel classes are generally within the same feature domain. Existing FSOD methods primarily rely on either fine-tuning or meta-learning paradigms [56]. Fine-tuning usually requires less GPU resources but have less generalization power. In contrast, methods based on meta-learning have more robust generalization because of the class-agnostic training process and usually make more accurate inference but demand more GPU resources for building task episodes [60]. Multi-modal few-shot object detection (MM-FSOD), as a new approach for FSOD, aims at utilizing extra text information to enhance visual feature representation and provide potential zero-shot inference ability (see Fig. 1). Broadly speaking, multi-modal machine learning involves training on diverse sources of information, which can range from vision, language, acoustic, tactile, and others. Multi-modal learning has shown significant success in problems such as visual question-answering [10, 49], robotics [3, 30], continual learning  [4], and medical image analysis [27]. Multi-modality is also central to modern Generative AI models such as GPT-4 [37] and Gemini [14] which can process and produce text and images to perform complex tasks.
Figure 1: Different FSOD tasks. The classic FSOD task (top) involves only visual information. MM-FSOD (middle) introduces a language modality, which can provide extra information to improve FSOD performance. Our proposed CDMM-FSOD task (bottom) is tailored for cross-domain scenarios and extends MM-FSOD to use richer text and semantics. Above, we show a cross-domain example where a model might train on images and text of common data (such as birds) and would need to generalize to detection of less common, substantially different data (such as electrical circuits).
Despite the success of multi-modal object detection (MM-OD) methods on few-shot task, the source and the target data may have a substantial domain gap in some practical detection problems. Thus, MM-OD methods may have degraded performance in such cross-domain, data scarce scenarios. In Fig. 2, we visualize our problem description, where we train and test existing object detection methods on data where the source and target domains differ. We propose the cross-domain multi-modal few-shot object detection (CDMM-FSOD) task, aiming to bridge the domain gap while transferring information from the source domain to the target domain. We hypothesize that using rich text within a sentence can be more effective than relying solely on simple semantics. This property may be particularly true when images in the target domain require technical language to describe, since most models are trained on data common in daily life. For example, when detecting product defects or analyzing medical images, experts could use a professional training manual to learn how to distinguish the appearance of the objects and how to label the image. However, for neural network models, how to utilize such rich text information in helping with these detection task is still unclear. We aim to investigate whether and how rich text can help address such cross-domain data gaps. In Fig. 1, we illustrate the difference of our proposed CDMM-FSOD compared to classic FSOD and MM-FSOD. Our rich text contain detailed and technical descriptions of the training data, and our novel data (the electrical circuit in this example) has prominent domain shift.
Figure 2: Performance results on 10-shot object detection on multiple cross-domain, few-shot datasets. We observe substantial cross-domain degradation for existing multi-modal object detection methods Next-chat [59], GOAT [55], and ViLD [15], as well as for a single-modal detection method, Meta-DETR [60]. In contrast, our proposed method has stronger performance on out-of-domain data.
We choose a meta-learning version of Detection Transformer (DETR) [5], i.e. Meta-DETR [60] as our baseline because of its Transformer-based structure and excellent modality adaption capacity. Unlike the classic multi-modal setting, we define a rich text description for each training category which may contain technical terminology. These sentences function as the linguistic support set for our meta-learning process, analogous to the role played by the image support set in terms of functionality. We design a meta-learning multi-modal aggregated feature module to fuse the vision and the language embedding and map them to a class-agnostic feature space as the meta support feature. Furthermore, to ensure the model obtains the detailed knowledge from the rich semantics, we propose a rich semantic rectify module to align the generated language embedding with the ground truth language embedding. In our experiments, we demonstrate consider performance improvement by using rich text. We further discuss and investigate the effect of text length on performance.
Our contributions include:
  • To the best of our knowledge, we are the first to propose and study the cross-domain problem for MM-FSOD.
  • We propose a meta-learning multi-modal aggregated feature module that utilizes rich text semantics to alleviate the performance degradation due to domain gaps.
  • We propose a rich semantic rectify module that learns to align the generated language embedding and ground truth language embedding, enabling the model to better understand the text.
  • Our experiments on cross-domain detection data demonstrate that our approach outperforms existing methods, suggesting that multi-modal learning, when combined with rich text, is effective for detection of few-shot, out-of-domain data.
2 Related Works
2.1 Few-shot Object Detection
The FSOD task is to train a model to accurately detect novel categories where only a few labeled samples are available. The dominant training paradigms of FSOD include methods based on fine-tuning and meta-learning. Fine-tuning methods often first pre-train a model on base data with sufficient training samples for each category, and then directly fine-tune the model on novel categories with few samples. Representative works include TFA [56], FSCE [50], FSRC [45], DeFRCN [38], etc. In meta-learning methods, i.e. learning to learn, instead of aiming at specific classes, a model is trained to learn a metric ability through tasks (i.e. episodes), so that it can utilize a Siamese network [28] to achieve accurate classification by measuring the similarity between support and query features; thus, it is expected to learn class-agnostic knowledge [6]. Each training episode in meta-learning methods builds a support set and query set; an $n$-way $k$-shot support set consists of $n$ categories with $k$ training instances for each. Such work includes Meta-FRCN [18] and Meta-DETR [60].
To avoid catastrophic forgetting of the base categories, researchers tune FSOD models on both base and novel data. However, while fine-tuning methods directly transfer the model on novel data in a class-specific manner, meta-learning methods tend to result in class-agnostic models, because they aim at learning a metric scheme between query features and support features. TFA [56] demonstrates that fine-tuning methods have potential to surpass meta-learning methods in detection performance. However, FCT [20] shows that meta-learning is still a powerful solution for FSOD, and that such methods are better at rapidly adapting to new tasks, due to their task-level feature representation.
The traditional backbone networks for FSOD include Faster R-CNN [42], YOLO [41], and ViT [8]. More recently, Detection Transformer (DETR) [5], a novel object detection network, is gaining increasing popularity for FSOD due to its competitive performance compared to classical frameworks; furthermore, its transformer-based architecture facilitates integration with NLP tasks. In our work, we thus follow the meta-learning paradigm provided by Meta-DETR [60].
2.2 Prompt Learning and Multi-modal Few-shot Object Detection
Multi-modal object detection (MM-OD) extends the classic object detection task to incorporate sources of information beyond images. We consider language as the extra modality, so the task is to train a model on image-text data. Most multi-modal object detection networks that benefit from prompt learning, utilize a pre-trained vision-language representation (e.g. CLIP [40] is a standard choice. This strategy has been broadly adopted in open-vocabulary detection tasks (OVD), i.e. the model could recognize novel classes with only a image-text pairs. Such a multi-modal representation ability greatly enhances the generalization of the base model, but usually utilizes relatively simple language, such as only using the category names or attributes [15]. The development of Large Language Models has significantly ushered the development of multi-modal object detection. For example, Next-chat [59] proposes a Q&A-based object detection-instance segmentation model. YOLO-World [7] builds a visual-linguistic interactive model based on the idea of region-text contrastive.
Multi-modal few-shot object detection (MM-FSOD) has not been broadly explored. Han et al. [17] propose a framework that utilize soft prompt tokens to realize low-shot object detection. However, these methods only use the simple image-level description or category name as the language label for the objects in an image. Such text information is insufficient when there exists a large domain gap when transferring knowledge from base classes to a novel dataset. Our work leverages this scheme to build a multi-modal network, using richer and more complex language.
2.3 Cross-domain Few-shot Object Detection
Cross-domain few-shot object detection (CD-FSOD) aims at solving FSOD when the source and target feature domains are substantially different. MoFSOD [29] proposes a multi-domain FSOD benchmark that consists of 10 datasets from different domains to simulate real-life situations. Their $k$-shot sampling is defined as $k$ images per class, and they demonstrate that effectively using the pre-training dataset plays a significant role in improving the model performance. CD-FSOD [57] proposes a cross-domain FSOD benchmark with balanced data per category (i.e. $k$ instances for each categories), making it better align with the general $K$-shot FSOD benchmark settings [56]. Furthermore, they propose a novel distillation-based baseline. Acrofod [12] design an adaptive optimization strategy in selecting proper augmented data. We attempt to alleviate the cross-domain FSOD degradation issue with the help of multi-modal feature alignment, and we mainly follow the benchmark provided by [12] to evaluate our model.
3 Proposed Methods
3.1 Preliminaries
In the classical Few-Shot Object Detection (FSOD) task, there are a set of base classes $C_{B}$ and novel classes $C_{N}$. While each class in $C_{B}$ has abundant training data, each class in $C_{N}$ only has a small number of instances. We assume there is no overlap among classes in $C_{B}$ and $C_{N}$. The training data ${\mathcal{D}_{\rm FSOD}=\{(x,y)_{i}\}_{i=1}^{N_{D}}}$ consists of $N_{D}$ pairs of RGB images $x$ and ground-truth labels $y$. Here, each $y=\{(c_{j},b_{j})\mid j\in\{1,2,\ldots,N_{\rm obj}\}\}$ contains, for each of the $N_{\rm obj}$ object instances in image $x$, its class $c_{j}\in C_{B}\cup C_{N}$ and bounding box $b_{j}$. The aim is to train a model on $\mathcal{D}_{\rm FSOD}$ such that, at test time, it can effectively detect objects belonging to the novel classes [26, 56].
The Multi-Modal Few-Shot Object Detection (MM-FSOD) task extends FSOD to incorporate text information. Formally, we have a dataset $\mathcal{D}_{\text{MM-FSOD}}=\{(x,\ell,y)_{i}\}_{i=1}^{N_{D}}$, where each data point now includes an additional set of category-specific text labels $\ell$ for each of the $N_{\rm obj}$ instances in image $x$. Formally, =$\ell=\{l_{j}\mid j\in\{1,2,\ldots,N_{\rm obj}\}\}$. Each text $l_{j}$ can take a variety of forms, and is typically either the category name or a template sentence that contains the name, such as: This is a picture of <category>. It is also possible to generate $\ell$ by using an external image-text model, such as CLIP [40], to generate sentences from the image [17]. Finally, the Cross-Domain Multi-Modal Few-Shot Object Detection (CDMM-FSOD) task we study in this paper is an extension of MM-FSOD when the classes in $C_{B}$ and $C_{N}$ are substantially different.
3.2 Rich Text
We propose to manually define a fixed, rich text description for each category $c\in C_{B}\cup C_{N}$ to form the text $\ell$ for the training data. We denote the set of text descriptions as $W^{C}=\{w^{1},w^{2},\ldots,w^{C}\}$, where $C=|C_{B}\cup C_{N}|$ is the number of categories. Each text description contains multiple tokens with variable length of $M$, noted as $w^{i}=\{w^{i}_{1},w^{i}_{2},\ldots,w^{i}_{M}\}$, where the first token $w^{i}_{1}=\langle s\rangle$ is the start of the sentence, and $w^{i}_{M}=\langle/s\rangle$ is the end of the sentence. To form the set $W^{C}$, we use two strategies: (i) we refer to the Wikipedia content of each category and define each category from the aspects of color, shape, attribute, material, etc. (ii) besides the aspects from the first strategy, we extend the sentence by adding the depiction of the image context while keep other text unchanged. For example, the original sentence “Motorbikes have two wheels, a motor, and a sturdy frame” will be extended as “Motorbikes have two wheels, a motor, and a sturdy frame, usually ride by a person.” In the revised sentence, we introduce a common visual relationship between two categories: motorbikes and persons. Please refer to the Appendix for our complete text descriptions for all classes.
Figure 3: The overall structure of our model. Our proposed multi-modal feature aggregation module and rich text rectify module are in red blocks, the detailed structures of them are in Fig. 4. The multi-modal feature aggregation module is response for the cross-modality feature embedding mix. The rich text rectify module is used to reinforce the cross-modality understanding of the model. We design this end-to-end model that takes a set of support and query images, as well as a group of rich category text as input for training, and outputs the object detection results of the query images.
3.3 Proposed Architecture
We use Meta-DETR [60] as our baseline network, which is built upon deformable DETR [64]. DETR is a transformer-based [53] object detection model. It has a ResNet [22] backbone to obtain the high-level image features; and then an encoder is applied to get the encoder memory embedding; followed by a decoder that could convert the memory embedding back to the object proposal. For the backbone part, we follow Meta-DETR [60] and use a weight-shared self attention module to build the meta-learning scheme for the query and support images. The classic DETR network includes an encoder and a decoder module, both consists of six multi-head attention layers. Meta-DETR build the meta-learning module only at the first multi-head attention layer of the encoder module. For the meat-learning module, the query and support images will be firstly sent to a shared ResNet feature extractor to get the basic image features. Then the query and support features will be passed through a multi-head attention module so that they are in the same feature space. The support sample feature will then serve as the $K$ and $V$ matrices while the query image feature is $Q$ of a single-head self-attention module. This single-head self-attention module is response for the class-specific feature matching among the support and query feature. Meanwhile, in order to align with the class-agnostic learning task setting of meta-learning, this single-head self-attention module will be applied on a parallel branch to map the class-specific ‘support’-‘query’ relationship to a class-agnostic ‘task embedding’-‘query’ relationship. Superficially, Meta-DETR designs an independent trainable class-agnostic feature prototype as $V^{\prime}$ for the parallel single-head self-attention branch. And the support and query sample feature will still serve as $K^{\prime}$ and $Q^{\prime}$ respectively. In this way, the single-head self-attention module will achieve an important feature mapping so that the original class-specific support sample feature could be converted to class-agnostic prototype feature. During the test time, the model will only use the class-agnostic prototype as support features. And then the aggregated class-agnostic prototype will be sent to the input of DETR to generate the final detection output.
Figure 4: The upper block is the detail of our meta-learning multi-modal aggregation module and the lower on is the rich semantic rectify module. We use different colors for different feature branches. Please note that the rich semantic rectify module is for training process only and thus will not participate in the inferencing process.
The overall architecture of our proposed method is shown in Fig. 3. We add the rich text label as the auxiliary support feature. Then we aggregate the image and language support feature to train the single-head self-attention meta-learning module. Here we note this module as meta-learning multi-modal feature aggregation module. Besides, in order to ensure that the model properly learn the knowledge from the rich text information, we adopt a rich text semantic rectify module to bidirectionally generate language feature, and then enforce it to align with the ground-truth language feature.
3.4 Multi-modal Feature Aggregation
Multi-modal feature aggregation (MM Aggre.) is the core module to fuse the multi-modal information, which is capable of aggregating multiple support categories. The detailed architecture of this module is visualized in Fig. 4. we use the existing CLIP pre-trained model to extract the primary image feature of the support/query set and the primary semantic feature of the rich text. Then the support image feature will be processed by a RoIAlign layer and an average pooling layer to get the instance-level category prototypes. This is a common operation in object detection because the image feature could not represent the instance feature; there usually exists multiple foreground objects in one image. Simultaneously, similar to prior object detection works, the background, as an extra hidden label, is also initialized as one of the category prototypes. And then these features will be mapped to the same feature space through a multi-head self-attention module (i.e. the first DETR encoder layer).
Then we concatenate and average the image and semantic features as the overall support feature (i.e. the support vision feature will be evenly concatenated with the corresponding language feature from the same category). Then we use a single-head self-attention module to implement (i) the ‘support’-‘query’ feature mapping and (ii) the class-agnostic ‘task embedding’-‘query’ encoding matching.
For the the ‘support’-‘query’ feature mapping, the aggregated multi-modal feature will serve as key ($K$) and value ($V$) matrix for the single-head self-attention module. Specifically, given $C$ support categories, for multi-modal support feature ($S\in\mathbb{R}^{(C+1)\times d}$) and query image feature ($Q\in\mathbb{R}^{HW\times d}$), the feature mapping coefficient will be:
$A=\text{softmax}\left(\frac{QS^{T}}{\sqrt{d}}\right)$ (1)
In a query image, only certain foreground objects have matched feature with the support instances, while the remaining part (background) should be ignored. Therefore, we should only pay attention to certain area of the query feature map, and weaken those areas that do not match with any given support prototype. To this end, we apply a sigmoid function for support prototype $\sigma(S)$ as Value matrix, and combine it with coefficient $A$ to build the attention mechanism. Consequently, the output of the ‘support’-‘query’ feature mapping process would be:
$Q_{1}=A\sigma(S)\odot Q,$ (2)
where $\odot$ represents element-wise product. The output $Q_{1}$ is a refined query feature map that keep only the ‘foreground’ feature.
For the ‘task embedding’-‘query’ encoding matching, a set of class-agnostic task prototype ($T\in\mathbb{R}^{(C+1)\times d}$) is defined to replace the support class prototype. In other words, the class-specific support class prototype will be mapped to the pre-defined task prototype, so that the query feature will be matched with the task prototypes for the prediction process. The task prototype $T$ would serve as the $V^{\prime}$ matrix for the same single head self-attention module as the ‘support’-‘query’ feature mapping. The initialization of $T$ is similar to sinusoidal position embedding in transformer. The corresponding output will be:
$Q_{2}=AT$ (3)
The overall output of the multi-modal aggregation module is the element-wise addition of $Q_{1}$ and $Q_{2}$.
3.5 Rich Text Semantic Rectify
The rich text semantic rectify module (Rich Text Rect.) is depicted in Fig. 4. Our idea is to use a transformer encoder structure  [1], noted by $f_{\theta}$ to generate the category-level language sequence features of the fused support-query samples, and then the generated language feature is expected to be aligned with its ground truth feature. Additionally, this language generation process is bidirectional, which means we generate the sentence in both forward and backward direction, so that we ensure the model have robust image-text understanding ability.
For a given category $c$, its rich text tokens undergo refinement through the multi-head self-attention module that shared with the vision branch to get its semantic prototype $l^{c}$, as shown in Fig. 4. Additionally, the support and query features are averaged to create a composite feature $p$, which serves as the $K$ and $V$ matrices for a multi-head self-attention module $A_{3}$, as shown in Fig. 4. In $A_{3}$, the refined text tokens act as the $Q$ matrix. Subsequently, it generates the predicted text sequence $\hat{W}$ of length $M$. This entire process is duplicated for both forward and backward prediction components, i.e. the model will predict $\hat{W}$ from both left to right and right to left direction so that it could understand the rich text through its bidirectional context. The loss function is:
$\mathcal{L}_{rect}=\frac{1}{2}\left[\sum_{i=1}^{C}\sum_{j=2}^{M}-\mathrm{log}f_{\theta}(\hat{w}_{ij}-w_{ij}|p)+\sum_{i=1}^{C}\sum_{j=1}^{M-1}-\mathrm{log}f_{\theta}(\hat{w}_{ij}-w_{ij}|p)\right]$ (4)
4 Experiments
4.1 Implementation Details
We conduct our experiments on hardware consisting of 8 3090 GPUs for parallel training and testing. We borrow hyper-parameters from DETR [5] by setting the initial learning rate as 0.001, and using a batch size of 4 for 1-shot scenarios and 1 for 5-shot and 10-shot scenarios. The DETR encoder and decoder both have 6 self-attention layers (the first layer of the encoder is embedded in our meta-learning multi-modal aggregation module). We follow Afham et al. [1] for the architecture of the image-text transformer decoder.
[…]
Method Shot Backbone Modality ArTaxOr DIOR UODD
1 5 10 1 5 10 1 5 10
Meta-RCNN $\circ$  [58] ResNet50 Single 2.8 8.5 14.0 7.8 17.7 20.6 3.6 8.8 11.2
TFA w/cos $\circ$ [56] ResNet50 Single 3.1 8.8 14.0 8.0 18.1 20.5 4.4 8.7 11.8
FSCE $\circ$ [51] ResNet50 Single 3.7 10.2 15.9 8.6 18.7 21.9 3.9 9.6 12.0
DeFRCN $\circ$ [39] ResNet50 Single 3.6 9.9 15.5 9.3 18.9 22.9 4.5 9.9 12.1
Distill-cdfsod $\circ$ [57] ResNet50 Single 5.1 12.5 18.1 10.5 19.1 26.5 5.9 12.2 14.5
[…]
Our Method w/ self-built text DETR-R101 Multi 14.8 48.1 61.0 13.7 26.3 31.3 5.5 11.9 17.4
Our Method w/ LLM text DETR-R101 Multi 15.1 48.7 61.4 14.3 26.9 31.4 6.9 12.5 17.5
4.2 Datasets
Following the standard evaluation on the CD-FSOD benchmark provided by Xiong et al. [57], we use ArTaxOr [9], DIOR [33] and UODD [25] as our target domain few-shot datasets. ArTaxOr contains 7 arthropods categories, DIOR consists of 20 satellite images categories, and UODD has 3 underwater creature categories. We also provide results for the CD-FSOD benchmark proposed by Lee et al. [29] in the Appendix.
Furthermore, we test on three other recently introduced CD-FSOD benchmarks that adopt balanced sampling: Clipart1k [24], DeepFish [44], and NEU-DET [48]. Since their corresponding data splits for CD-FSOD are not officially published, we put the results in the Appendix.
To measure the generalization of our method, we evaluate on the general FSOD benchmark of PASCAL VOC. Following the pioneering evaluation pipeline for FSOD [56], 15 of the categories of PASCAL VOC are selected as base classes for pre-training, and the remaining 5 categories are used as the novel classes for few-shot fine-tuning. See the Appendix for a detailed description.
4.3 Results on CD-FSOD Datasets
Our experimental results on CD-FSOD datasets are presented in Tab. 1. We follow the experimental setting from Xiong et al. [57]; our model is pre-trained on COCO and fine-tuned on the three out-of-domain datasets. We report the average precision (mAP) for evaluation. Since there is no existing work on CDMM-FSOD, we implement representative state of the art FSOD and MMOD works on the CD-FSOD benchmark and compare our method with them. The values reported by Distill-cdfsod [57] and Fu et al. [11] are directly from the papers; while the values of other methods (i.e. Meta-DETR, Next-Chat, GOAT and ViLD) are generated by us. We run the official code of these methods on the CD-FSOD benchmark. The experimental results indicate that our method outperforms the baselines. It is worth noting that we achieve significant improvement on the ArTaxOr dataset. Specifically, with LLM text, we reach 15.1, 48.7, 61.4 for 1,5,10-shot scenarios compared to the highest baseline values of 6.5, 20.9, 23.4, respectively. The reason is that the 7 categories of ArTaxOr have less domain shift compared to the COCO pre-training dataset in terms of the more clearer object boundary compared to other two cross-domain benchmarks [11].
As introduced in Sec. 3.2, we build rich text manually or by using an LLM [52]. As indicated in Tab. 1, either way could generate significant improvement on the out-of-domain benchmarks. It is observed that LLM-generated rich text brings better performance on these three benchmarks, because the generated text contains more comprehensive sentences. However, this is not a universal situation, because sometimes the LLM failed to generate distinct descriptions for the target categories when their names are abstract and technical, e.g. in NEU-DET dataset; for further discussion please refer to the Appendix.
4.4 Results on Standard FSOD Datasets
To further evaluate the generalization of our method, we test our model on a common FSOD benchmark, i.e. few-shot PASCAL VOC, in Tab. 2. Following TFA [56], as described in Sec. 4.2, there are three fixed groups of base-novel category splits, namely split 1, 2, and 3 to avoid sampling bias. For each data split, the $n$-shot ($n\in\{1,2,3,5,10\}$) sampling for the novel categories are also pre-defined. The model is evaluated on these three respective dataset splits. In addition, as described in Sec. 3.1, the base categories are also trained during fine-tuning to avoid forgetting, but have different sampling strategies: (i) in works such as TFA [56], the base categories also follow the restricted $n$-shot sampling (i.e. $n$ instances instead of $n$ images for each categories) as the novel categories; (ii) while in works such as Meta-DETR [60], the base categories are not few-shot. Both strategies are acceptable for the FSOD task; we adapt to the later one. As shown in Tab. 2, our method outperforms previous works.
Table 2: Performance results on PASCAL VOC dataset w/ self-built rich text descriptions. Highest scores are bolded.
Method Shot Backbone Split1 Split2 Split3
1 2 3 5 10 1 2 3 5 10 1 2 3 5 10
TIP [31] FRCN-R101 27.7 36.5 43.3 50.2 59.6 22.7 30.1 33.8 40.9 46.9 21.7 30.6 38.1 44.5 50.9
CME [32] FRCN-R101 41.5 47.5 50.4 58.2 60.9 27.2 30.2 41.4 42.5 46.8 34.3 39.6 45.1 48.3 51.5
DC-Net [23] FRCN-R101 33.9 37.4 43.7 51.1 59.6 23.2 24.8 30.6 36.7 46.6 32.3 34.9 39.7 42.6 50.7
CGDP [35] FRCN-R2101 40.7 45.1 46.5 57.4 62.4 27.3 31.4 40.8 42.7 46.3 31.2 36.4 43.7 50.1 55.6
Meta Faster R-CNN [19] FRCN-R101 40.2 30.5 33.3 42.3 46.9 26.8 32.0 39.0 37.7 37.4 34.0 32.5 34.4 42.7 44.3
FCT [21] PVTv2-B2-Li 49.9 57.1 57.9 63.2 67.1 27.6 34.5 43.7 49.2 51.2 39.5 54.7 52.3 57.0 58.7
Meta-DETR [60] DETR-R101 40.6 51.4 58.0 59.2 63.6 37.0 36.6 43.7 49.1 54.6 41.6 45.9 52.7 58.9 60.6
MM-FSOD [17] ViT 42.5 41.2 41.6 48.0 53.4 30.5 34.0 39.3 36.8 37.6 39.9 37.0 38.2 42.5 45.6
Our Method DETR-R101 45.1 54.9 62.2 65.1 69.6 41.7 42.1 47.7 54.3 57.3 49.2 52.2 56.9 63.1 64.0
4.5 Ablation Study
We first verify that the length of the rich text sentences have a distinct impact on the model’s performance. To test ‘length’, we conduct experiments on the UODD dataset using 4 cases of text: (i) the raw category name, (ii) our manually-built rich text referencing Wikipedia, (iii) an extended version of (ii) by adding image context, (iv) LLM-generated rich text. Cases (ii) to (iv) all result in long sentences that have rich semantics, and thus are generally beneficial for the model performance. See Tab. 3 for our examples and quantitative results. However, not all extensions on the sentence length are effective. For example, in case (iii), our extended rich text for sea cucumbers has a slightly negative impact over the non-extended version, decreasing performance from 17.4 to 16.9. This may be because the extra image context with an distinct name of another category in UODD (i.e. sea urchin) has confused the model.
In addition, we analyze the effectiveness of our proposed multi-modal aggregated feature module and the rich semantic rectify module. As seen in Tab. 4, comparing to the single-modal baseline, our method can achieve significant improvement through using rich text multi-modal information. Our rich text rectify module also leads to an additional improvement.
Table 3: The effect of different language modalities.
Type of text
Example
UODD
None
e.g. None
9.2
Category name
e.g. Sea cucumbers
9.9
Rich text
e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles
17.4
Extended rich text
e.g. Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles; usually seen together with sea urchins.
16.9
LLM-Rich text
e.g. Sea cucumbers are marine invertebrates known for their elongated, leathery bodies, which are typically covered in spines or tentacles and lack a distinct head or tail
17.5
Table 4: The effect of different modules.
Module ArTaxOr DIOR UODD
Meta-DETR 14.5 19.9 9.2
Meta-DETR + MM Aggre. 59.8 30.1 15.7
Meta-DETR + MM Aggre. + Rich Text Rect. 61.4 31.4 17.5
Figure 5: Representative visualizations of detection results on three benchmark datasets. We compare our method with a multi-modal object detection model (Next-chat) and a few-shot object detection model (Meta-DETR). Our proposed model obtains more accurate bounding boxes and improved detection confidence.
4.6 Visualization
In Fig. 5, we present representative qualitative detection results on the three CD-FSOD benchmarks. In these examples, the objective is to detect araneae (ArTaxOr), sea urchins and sea cucumbers (UODD), and windmills (DIOR). The inference uses a confidence threshold setting of 0.3 for all methods. Compared to baseline methods, we show qualitative improvements in a more accurate bounding box (first row), fewer false positives (second row), and improved detection confidence (third row). See the Appendix for additional qualitative results.
4.7 Discussion
We observe that it is desirable when the rich text description of the target objects are available, but the specific context of the rich text plays a critical role. Consequently, in practice, the user must carefully evaluate the rich text and substantially tune it for greater improvement. This might become a burden in some cases when (i) the user has a rich text, but it contains many descriptions that are irrelevant to the visual information, or (ii) the user does not have rich text and thus need to build one from scratch. For the first case, a general suggestion for practical usage is to keep the information most related to the appearance of the target objects. For the second case, a user could extend and improve upon our LLM-generated text approach. In future work, for example, we plan to train an discriminator to select the best rich text from multiple candidates.
5 Conclusion
We introduce a novel cross-domain multi-modal few-shot object detection network. We verify that using rich text information is a promising strategy for robust detection in out-of-domain, few-shot scenarios. Our experiments indicate that the design of rich text is a key impact factor for the model’s performance, and that a precise description of the appearance of the target objects is especially effective. Our experimental results demonstrate improved precision over current state of the art methods on multiple cross-domain object detection datasets. We hope that this paper inspires future work to explore using multi-modality for bridging domain gaps in other computer vision tasks.
Appendix 0.A Experiments on Additional Benchmarks
In this section, we present our results on other CD-FSOD benchmarks.
Fu et al. [11] introduce one such benchmark built on three datasets: Clipart1k [24], DeepFish [44], and NEU-DET [48]. The few-shot sampling strategy is balanced (i.e. $k$ shot refers to $k$ instances). The results are presented in Tab A5, demonstrating that our approach outperforms previous methods, particularly on the Clipart1k dataset. Specifically, in the 10-shot scenario, we achieve a score of 47.7 and 48.8, significantly surpassing the best prior method’s score of 25.6. However, we observe that in the NEU-DET dataset, the text generated with the assistance of LLM does not yield superior results compared to manually crafted text. This is because the LLM failed to produce texts with sufficient distinguishing features as seen in other datasets (see Section 0.E.4). Specifically, the LLM-aided text descriptions for the six defect categories tended to exhibit homogeneity, limiting their effectiveness.
Lee et al. [29] introduce another CD-FSOD benchmark which includes 10 datasets: VisDrone [63] for aerial images, DeepFruits [43] for agriculture, iWildCam [2] for animals in the wild, Clipart [24] for cartoons, iMaterialist [16] for fashion, Oktoberfest [65] for food, LogoDet-3K [54] for logos, CrowdHuman [47] for people, SIXray [36] for security, and KITTI [13] for traffic/autonomous driving. However, these benchmarks use an unbalanced sampling strategy, i.e., $k$-shot means $k$ images per class. We conduct experiments on two most visual unique datasets: Clipart and SIXray, as illustrated in Tab. A6. Our method surpass the performance of Lee’s result.
[…]
Method Shot Backbone Modality Clipart1k DeepFish NEU-DET
1 5 10 1 5 10 1 5 10
[…]
Our Method w/ self-built text DETR-R101 Multi 22.9 42.6 47.7 21.2 23.3 25.2 6.1 15.2 19.9
Our Method w/ LLM text DETR-R101 Multi 25.6 44.3 48.8 23.5 25.6 26.4 5.8 13.1 13.2
Table A6: Performance Results (mAP) on CD-FSOD benchmarks proposed by Lee et al. [29]. Highest scores are in bold font.
Method Shot Backbone Modality Clipart SIXray
1 5 1 5
Lee et al. ResNet-50 Single 37.2 49.3 6.6 23.9
Our Method w/ self-built text DETR-VIT-L Multi 46.5 58.4 13.3 34.7
Our Method w/ LLM text DETR-VIT-L Multi 47.2 58.8 14.5 35.3
Appendix 0.B General FSOD Benchmark
The general FSOD framework enables the model to undergo fine-tuning on both novel and base data, helping to mitigate catastrophic forgetting. The benchmark evaluation for general FSOD, as proposed by TFA [56], on the PASCAL VOC dataset comprises 20 categories, with 15 designated as base categories and the remaining 5 as novel categories. To eliminate any potential data bias, 3 fixed category splits have been established for training and testing, enabling an accurate assessment of the model’s average performance. The benchmark employs two distinct sampling strategies for sampling the base categories during fine-tuning: balanced sampling and unbalanced sampling. In the case of balanced sampling, the $k$-shot refers to utilizing $k$ instances for each category, including both novel and base categories. This sampling approach is employed in methods like TFA. On the other hand, the unbalanced sampling strategy involves the $k$-shot specifically referring to strictly selecting $k$ instances for the novel categories alone, excluding the base categories. This sampling approach is employed in methods like Meta-DETR [60]. Both strategies are acceptable for general FSOD, and for our results reported in Tab. 2, we follow the unbalanced sampling strategy.
Appendix 0.C Visualization of the Detection Result
We present additional visualizations of our detection results as an extension of Sec. 4.6. The detection results on the ArTaxOr, DIOR and UODD dataset are illustrated in Fig. A6, Fig. A7 and Fig. A8 respectively. We demonstrate consistent improvements compared to the multi-modal method (Next-chat) and the single-modal method (Meta-DETR). Specifically, in Fig. A6, we observe more precise classification of Lepidoptera, improved bounding box regression for Odonate, and enhanced classification confidence. Similarly, notable advancements are also evident in Fig. A7 and Fig. A8. Furthermore, in Fig. A8, we observe additional foreground detection results, such as the sea cucumber in the second row and the scallop in the third row.
Figure A6: Examples of detection results from ArTaxOr.
Figure A7: Examples of detection results from DIOR.
Figure A8: Examples of detection results from UODD.
Appendix 0.D Computational Overhead
We list our computational overhead as a reference in Tab. A7. The values are reported based on 10-shot (68 images each epoch) experiments on the ArTaxOr dataset. See Sec. 4.1 for additional hyperparameters we use. Our method has more than 10X training parameters compared to Meta-DETR, but only has about 2X training time because we calculate and cache the feature embedding of the rich text.
Table A7: Computation overhead of our methods
Trainable parameters Model size Training time
Meta-DETR 11,649,794 317MB 36 s/epoch
Our method 162,927,107 671MB 84 s/epoch
Appendix 0.E Detailed List of the Rich Text
0.E.1 ArTaxOr Dataset
As depicted in Fig. A9, the ArTaxOr Dataset [9] comprises seven distinct categories of insects. The images within this dataset were primarily captured using a macro lens, resulting in high-definition visuals and a distinct foreground-background boundary. Nevertheless, the inter-class differences within the ArTaxOr Dataset are relatively subtle, posing a challenge in accurately distinguishing between the various insect categories [57].
Figure A9: Examples of images and categories from ArTaxOr.
  • Araneae:
    • (manually built) Araneae have eight limbs and usually perch on the silk, and do not have antennae.
    • (LLM-aided) Araneae, commonly known as spiders, exhibit a diverse range of appearances, varying from small and delicate to large and imposing, with eight legs, two body segments, and often distinctive markings and patterns on their abdomens.
  • Coleoptera:
    • (manually built) Beetles can usually be recognized by their two pairs of wings; the front pair is modified into horny covers that hide the rear pair and most of the abdomen and usually meet down the back in a straight line.
    • (LLM-aided) Coleoptera, the beetles, exhibit a remarkable diversity in appearance, ranging from small, shiny, and flightless to large, winged, and brightly colored, often with hardened forewings adapted for protection or flight.
  • Diptera:
    • (manually built) Diptera use only a single pair of wings to fly, and have a mobile head, with a pair of large compound eyes, and mouthparts designed for piercing and sucking.
    • (LLM-aided) Diptera, or flies, are small insects with two wings, large compound eyes, and long antennae, varying in color from dull to bright, with some species having patterns or stripes on their bodies.
  • Hemiptera:
    • (manually built) Hemiptera’s forewings are hardened near the base and membranous near the ends.
    • (LLM-aided) Hemiptera, or bugs, are small to medium-sized insects typically characterized by a flattened body, strong forewings hardened into protective plates, and prominent eyes. Some species are brightly colored while others are dull-brown or black.
  • Hymenoptera:
    • (manually built) Hymenoptera usually have two pairs of membranous wings and the forewings are larger than the hind wings.
    • (LLM-aided) Hymenoptera are distinguished by their thin waist connecting the thorax and abdomen, as well as their typically transparent and membranous wings.
  • Lepidoptera:
    • (manually built) Lepidoptera have scales that cover the bodies, large triangular wings, and a proboscis for siphoning nectars.
    • (LLM-aided) Lepidoptera, including butterflies and moths, are characterized by their wings, which are covered in scales and often display patterns and colors ranging from subtle to vivid. Their bodies are typically slender, and they have antennae and long, thin legs.
  • Odonata:
    • (manually built) Odonata characteristically have large rounded heads covered mostly by compound eyes, two pairs of long, transparent wings that move independently, elongated abdomens, three ocelli and short antennae.
    • (LLM-aided) Odonata, commonly known as dragonflies and damselflies, are characterized by their large, often brightly colored eyes, long, thin bodies, and two pairs of wings, the hindwings being broader than the forewings.
0.E.2 UODD Dataset
The UODD dataset [25], as illustrated in Fig. A10, comprises three distinct categories of underwater creatures. Due to the low visual contrast, it exhibits significant differences from the COCO dataset [57].
Figure A10: Examples of images and categories from UODD.
  • Sea cucumber:
    • (manually built) Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles.
    • (extended) Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles; usually seen together with sea urchins.
    • (LLM-aided) Sea cucumbers are marine invertebrates known for their elongated, leathery bodies, which are typically covered in spines or tentacles and lack a distinct head or tail.
  • Sea urchin:
    • (manually built) Sea urchins have globular body and a radial arrangement of organs.
    • (extended) Sea urchins have globular body and a radial arrangement of organs; usually seen together with sea cucumbers.
    • (LLM-aided) Sea urchins are marine invertebrates characterized by their globular shape, covered in dense clusters of spines for protection, and having a distinctive oral surface with five radiating rows of teeth-like structures called pedicellariae.
  • Scallop:
    • (manually built) The shell of a scallop has the classic fanned-out shape; one side of the shell is slightly flatter, and the other more concave in shape.
    • (extended) Sea cucumbers have sausage-shape, usually resemble caterpillars; their mouth is surrounded by tentacles; usually seen together with sea urchins.
    • (LLM-aided) Scallops are marine bivalve mollusks characterized by their flattened, oval-shaped shells, typically with distinctive radial ribbing and a wavy or scalloped edge.
0.E.3 DIOR Dataset
As depicted in Fig. A11, the DIOR dataset [33] encompasses 20 categories of Aerial scenes. These scenes differ significantly from those in the COCO dataset due to their unique remote sensing perspective, captured from a top-down shot [57].
Figure A11: Examples of images and categories from DIOR.
  • Expressway service area:
    • (manually built) A expressway service area is an open space located alongside the highway, typically featuring a main building and multiple regularly arranged parking slots.
    • (LLM-aided) Expressway service areas appear as clusters of buildings and facilities along the highway, typically surrounded by parking lots and green spaces.
  • Expressway toll station:
    • (manually built) Toll stations on expressways are elongate, thin buildings that span across the highway, and allow vehicles to queue up for entry or exit.
    • (LLM-aided) Expressway toll stations appear as structures located along the highway, typically with multiple lanes entry and exit points, and are distinguished by their regular layout, clear lanes markings, and sometimes by the presence of toll plazas or gantries.
  • Airplane:
    • (manually built) An airplane is equipped with a pair of symmetric, large wings positioned at the center of its cylindrical fuselage, along with a pair of smaller wings referred to as the horizontal stabilizer and a vertical stabilizer located at the tail end.
    • (LLM-aided) Airplanes are winged vehicles with a fuselage, wings, and engines, typically having a pointed nose and tail.
  • Airports:
    • (manually built) Airports are designated open areas where airplanes can park, and they typically feature multiple comb-shaped boarding bridges, tall control towers, and lawns, but lack trees.
    • (LLM-aided) Airports are large open spaces with multiple runways, terminals, and parking lots, typically surrounded by fences and gates, and equipped with aircrafts, control towers.
  • Baseball field:
    • (manually built) The baseball field features a diamond-shaped infield, resembling a quarter of a circle, covered with a checkerboard pattern of natural green grass.
    • (LLM-aided) A baseball field is characterized by a diamond-shaped infield with four bases, a pitcher’s mound, and outfield fences, surrounded by stands for spectators and often marked by lines indicating playing areas and distances.
  • Basketball court:
    • (manually built) The basketball court is the playing surface, consisting of a rectangular floor with tiles at either end, usually made out of a wood, often maple, and highly polished.
    • (LLM-aided) A basketball court is a rectangular playing surface with a raised rim and net at each end, typically surrounded by lines dividing the court into various zones and marked by a center circle and free-throw lines.
  • Bridge:
    • (manually built) Bridges are elongate, slender structures spanning across water bodies and may accommodate the passage of vehicles.
    • (LLM-aided) Bridges are structures built to span rivers, valleys, or other obstacles, typically with one or more arches, trusses, or cables, connecting two or more points of land.
  • Chimney:
    • (manually built) A chimney is like a pipe or a tunnel-like channel made by metal or concrete, typically tall and expel gas.
    • (LLM-aided) Chimneys are vertical structures designed to vent smoke and gases from heating appliances, typically made of masonry or metal and having a rectangular or circular cross-section.
  • Dam:
    • (manually built) A dam is a structure constructed across a stream, river, or estuary to retain water, and from an overhead perspective, it typically appears as a thin, curved white line.
    • (LLM-aided) Dams are large structures built across rivers or streams to control water flow, typically with a wall or gates to hold back water and create a reservoir, and may include spillways to control flooding.
  • Golf course:
    • (manually built) Golf courses are extensive, irregularly shaped open spaces characterized by meticulously manicured lawns, and the terrain is typically gently undulating.
    • (LLM-aided) Golf courses appear as large, open spaces with regularly spaced tee boxes, greens, and fairways, typically surrounded by fences or hedges and distinguished by their smooth, manicured surfaces.
  • Ground track field:
    • (manually built) Ground track fields typically consist of red rings forming elliptic paths, with a pair of straight lines facing each other and another pair of half-circles extending in the opposite direction.
    • (LLM-aided) Ground track fields appear as large, open spaces with straight lines and rectangles marking the field boundaries and flight paths, typically surrounded by fences or other markers and distinguished by their regularity and symmetry.
  • Harbor:
    • (manually built) A harbor refers to a sheltered body of water, either natural or man-made, and it is commonly seen filled with boats or ships parked within its confines.
    • (LLM-aided) Harbors appear as clusters of ships and boats moored along the coast or riverbanks, typically surrounded by docks, piers, and other facilities for loading and unloading cargo, and distinguished by their dense concentration of vessels and associated infrastructure.
  • Overpass:
    • (manually built) Overpasses are intricate city roadways that resemble bridges and intersect each other, typically elevated above standard road levels to facilitate the passage of vehicles.
    • (LLM-aided) Overpasses appear as elevated structures spanning roads or railroads, typically with multiple arches or girders supporting the roadway above the traffic below, and distinguished by their raised position and distinctive shape.
  • Ship:
    • (manually built) Ships typically feature a long, thin, bullet-shaped hull that floats on the water, with a sharp prow and a flat stern.
    • (LLM-aided) Ships appear as large, elongated vessels floating on the water surface, typically with a distinctive silhouette and various features such as decks, masts.
  • Stadium:
    • (manually built) Stadiums resemble bowl-shaped buildings and typically feature stands surrounding a central field.
    • (LLM-aided) Stadiums appear as large, open spaces with distinctive curved or rectangular shapes, typically surrounded by stands and other facilities for spectators.
  • Storage tank:
    • (manually built) Storage tanks are usually large cylindrical or spherical metal containers with a clean exterior, clustered and well-arranged on the ground.
    • (LLM-aided) Storage tanks appear as large, circular or rectangular structures with flat roofs, typically surrounded by fenced areas and distinguished by their regular shape, smooth surfaces, and sometimes by the presence of multiple tanks clustered together.
  • Tennis court:
    • (manually built) Tennis courts are rectangular fields marked with straight lines across their surfaces, and they are typically colored green or blue.
    • (LLM-aided) Tennis courts appear as rectangular areas with distinctive lines marking the boundaries and service lines, typically surrounded by fences or hedges and distinguished by their regularity and symmetry.
  • Train station:
    • (manually built) Train stations are large buildings designed to accommodate trains, and they typically feature a dense network of multiple railways passing through them.
    • (LLM-aided) Train stations appear as clusters of buildings and platforms along the railway tracks, typically with multiple tracks converging and diverging, and distinguished by the presence of trains or other rail vehicles.
  • Vehicle:
    • (manually built) Vehicles are typically small rectangular objects commonly seen traveling along roads or parked in parking lots.
    • (LLM-aided) Vehicles appear as small to medium-sized objects with distinctive shapes and features, such as wheels, windows, and roofs, typically moving along roads or parking areas.
  • Windmill:
    • (manually built) Windmills are tall structures topped with multiple fan-like blades that spread out and rotate to harness wind energy.
    • (LLM-aided) Windmills appear as tall, thin structures with three or more long blades rotating in the wind, typically located in open spaces such as fields or coastal areas.
0.E.4 NEU-DET Dataset
As shown in Fig. A12, the NEU-DET dataset [48] features six categories of common steel rolling product defects. These images were captured using an electron microscope, resulting in a distinct black-and-white style. Notably, the industrial image style of the NEU-DET dataset differs significantly from the life scenes presented in the COCO datasets.
Figure A12: Examples of images and categories from NEU-DET.
  • Rolled-in scale:
    • (manually built) Rolled-in scale typically consists of multiple rough patches arranged in a fish scale-like pattern.
    • (LLM-aided) In hot-rolled steel strip defect images, rolled-in scale appears as dark, irregular patches or streaks along the surface of the steel strip.
  • Patches:
    • (manually built) Patch defects typically appear as dark, dense clusters.
    • (LLM-aided) In hot-rolled steel strip defect images, a patch typically appears as a localized area of abnormal texture, color, or brightness
  • Crazing:
    • (manually built) Crazings typically consist of numerous tiny, thin lines resembling tree branches.
    • (LLM-aided) Crazing in hot-rolled steel strip defect images is characterized by a network of fine cracks or fissures that appear on the surface of the steel strip, typically have a regular, hairline pattern.
  • Pitted surface:
    • (manually built) Pitted surfaces are characterized by a dense cluster of tiny spots.
    • (LLM-aided) Pitted surfaces are characterized by numerous small, circular indentations or depressions spread across the steel strip’s surface.
  • Inclusion:
    • (manually built) Inclusions typically consist of a series of dark, slender patches.
    • (LLM-aided) Inclusions in hot-rolled steel strip defect images appear as dark, irregularly shaped areas within the steel strip’s matrix, typically caused by foreign particles or impurities during the manufacturing.
  • Scratches:
    • (manually built) Scratches are typically characterized by multiple light, straight, slender lines.
    • (LLM-aided) Scratches appear as long, continuous lines or marks across the steel strip’s surface.
0.E.5 Clipart1k Dataset
The Clipart1k Dataset [24], in Fig. A13, comprises 20 categories of commonly seen objects that overlap with the COCO categories, however, due to their cartoon-like style, models often struggle to accurately recognize them.
Figure A13: Examples of images and categories from Clipart1k.
  • Aeroplane:
    • (manually built) The cartoon aeroplane typically features a pair of wings positioned in the center of its cylindrical body, accompanied by a set of small stabilizers at the rear, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon aeroplane is characterized by its sleek fuselage, wings and various mechanical components, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Bicycle:
    • (manually built) The cartoon bicycle typically comprises of two wheels, one at the front and another at the rear, along with a triangular-shaped frame, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon bicycle is typically defined by its two wheels, a frame connecting them, handlebars for steering, and a seat for the rider, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Bird:
    • (manually built) The cartoon bird commonly has a furry body, a pair of feathery wings, and a pair of slender feet, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon bird is typically characterized by its feathered body, beaked head, and wings, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Boat:
    • (manually built) The cartoon boat typically features a half-moon-shaped hull floating on the water, and sometimes with a set of big sails, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon boat is typically defined by its hull, which floats on water, along with features such as a deck, mast, and sails or engines, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Bottle:
    • (manually built) The cartoon bottle commonly exhibits a transparent cylindrical body, characterized by a slender neck and a wider abdomen, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon bottle is typically characterized by its distinct shape, often cylindrical or tapered, with a narrow neck and a mouth for pouring or sealing, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Bus:
    • (manually built) The cartoon bus typically takes the form of a cuboid-shaped vehicle, supported by four wheels at the base. Its body is surrounded by rectangular windows, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon bus is typically characterized by its large size, rectangular shape, and multiple doors for passengers, along with windows for visibility and a roof that often features an advertising panel or other decorative elements, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Car:
    • (manually built) Cartoon cars are vehicles typically featuring a small, box-shaped driving cab and an flat elongated front engine box, supported by four wheels at the base, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon car is typically defined by its sleek and compact body, with four wheels, a windshield, and various features such as headlights, taillights, and mirrors, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Cat:
    • (manually built) Cartoon cats possess a furry body with four limbs, a pair of triangular ears on top of their heads, long tails, and slightly pointed mouths, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon cat is typically characterized by its sleek and muscular body, furry coat, pointed ears, and large, expressive eyes, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Chair:
    • (manually built) The cartoon chair typically features four slender legs for support, along with a small, horizontal, flat area designed for sitting. sometimes includes a vertical backrest and a pair of armrests, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon chair is typically defined by its supportive structure, consisting of a seat, backrest, and legs, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Cow:
    • (manually built) The cartoon cow commonly possesses four long legs that support its box-shaped body, with a long tail, sometimes, it may also feature a pair of crescent-shaped horns protruding from its head, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon cow is typically characterized by its large, stocky body, brown or black fur, and prominent horns, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Diningtable:
    • (manually built) The cartoon dining table is typically a large, tall piece of furniture supported by multiple legs; it’s surface is often rounded or rectangular, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon dining table is typically defined by its rectangular or circular shape, flat surface for placing food and dishes, and supportive legs, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Dog:
    • (manually built) The cartoon dog typically exhibits a furry body, characterized by a prominent, elongated nose and mouth, leaf-shaped ears, four legs, and a tail that is often curled, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon dog is typically characterized by its varied coats, distinct ears and muzzle shapes, as well as its four legs and tail, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Horse:
    • (manually built) The cartoon horse typically stands on four legs, supporting its body, and features a brush-like tail, a long neck and a long face, as well as a long mane along its neck, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon horse is typically characterized by its powerful legs, long mane and tail, and muscular body, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Motorbike:
    • (manually built) The cartoon motorbike typically features two bold wheels at the front and rear of its body, and a prominent engine positioned in the middle, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon motorbike is typically defined by its compact and aerodynamic frame, two wheels, a handlebar for steering, and an engine for power, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Person:
    • (manually built) The cartoon person typically possesses two legs to support their body, a pair of arms, and a head with a pair of eyes, ears, and a mouth, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon person is typically characterized by a bipedal stance, upright posture, and a head with facial features such as eyes, nose, and mouth, along with arms, legs, and varying skin tones and hairstyles, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Potted plant:
    • (manually built) The cartoon potted plants are typically small trees or bushes grown indoors in pots, with green leaves or vibrant flowers, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon potted plant is typically defined by its green foliage or flowers, growing from a soil-filled pot, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Sheep:
    • (manually built) The cartoon sheep commonly stands on four legs, supporting its body, and is covered in white, curly fur; it features a short, curly tail and a pair of horns that curl downward, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon sheep is typically characterized by its fleecy white coat, rounded body, and curved horns, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Sofa:
    • (manually built) The cartoon sofa boasts a long, box-shaped body designed for seating, adorned with cushions and featuring a thick backrest as well as a pair of sturdy armrests, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon sofa is typically defined by its cushioned seating, backrest, and arms, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • Train:
    • (manually built) The cartoon train comprises a succession of long, box-shaped compartments that traverse along a rail, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon train is typically characterized by its long and connected series of carriages, each with windows and doors, a powerful locomotive at the front, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
  • TV monitor:
    • (manually built) The cartoon TV monitor typically comprises a screen for displaying images and a sturdy base for support, usually in unified colors and clear contours.
    • (LLM-aided) A cartoon TV monitor is typically defined by its rectangular screen, displaying images and sounds, with a frame or border, and various buttons or controls for operation, all exaggerated in features, enhanced with brighter colors, and adorned in a whimsical or anthropomorphic design.
0.E.6 DeepFish Dataset
The DeepFish dataset[44], as illustrated in Fig. A14, exclusively features a single category of high-definition underwater tropical fishes, all sampled in Australia. Its unique image style, characterized by blurred visual boundaries and the fishes’ biological camouflage, sets it apart from the COCO dataset.
Figure A14: Examples of images and categories from DeepFish.
  • Fish:
    • (manually built) Underwater fish typically exhibit a spindle-shaped body covered with scales, along with a pair of pectoral fins and a tail that is crescent-shaped.
    • (LLM-aided) Underwater fish typically exhibit diverse appearance features, ranging from sleek and streamlined bodies to brightly colored scales and fins.
0.E.7 SIXray Dataset
The SIXray dataset[36], as illustrated in Fig. A14, exclusively features six category of contraband for security check. Its unique X-ray image style, characterized by complex boundaries, sets it apart from the COCO dataset.
Figure A15: Examples of images and categories from SIXray.
  • Knife:
    • (manually built) Knives appear as long, dark shadows, consisting of a handle and a sharp blade.
    • (LLM-aided) The knife typically exhibits a blade with a sharp edge for cutting
  • Gun:
    • (manually built) Guns are discernible as dark L-shaped shadows, featuring a long, slender barrel on one side and a shorter, stubby handle on the other.
    • (LLM-aided) The gun typically exhibits a compact and handgun-like shape, featuring a barrel, a grip for holding, a trigger for firing.
  • Wrench:
    • (manually built) The wrench is discernible as a slim, elongated shape, with one end featuring a horseshoe-shaped head and the other end attached with a circular ring.
    • (LLM-aided) The wrench typically features a handle for gripping and a jaw or socket that can be adjusted to fit nuts, bolts, or other fasteners, enabling it to rotate or tighten them.
  • Pliers:
    • (manually built) Pliers are characterized by their two slender handles, connected to a pointed, snip-nosed jaw.
    • (LLM-aided) The pliers typically consist of two pivoting levers with jaws at the ends, designed to grasp and manipulate objects, with handles for easy gripping and control.
  • Scissors:
    • (manually built) Scissors comprise two circular handles connected to a long, sharp blade that forms the cutting jaw.
    • (LLM-aided) The scissors typically consist of two pivoting blades with sharp edges that meet at a point, allowing for cutting action when the handles are squeezed together.